# AI-Powered Inverse Design of Ku-Band SIW Resonant Structures by Iterative Residual Correction Network

### v1

@@ Abstract
Inverse electromagnetic modeling has emerged as a powerful approach for designing complex microwave structures with high accuracy and efficiency. In this study, we propose an Iterative Residual Correction Network (IRC-Net) for the inverse design of Ku-band Substrate Integrated Waveguide (SIW) components based on multimode resonators. We use a multimode resonance structure to demonstrate that it is possible to control the resonances of the structure. Therefore, these structures can be used for resonant components and smart filter design. The proposed deep learning architecture leverages residual neural networks to overcome the limitations of traditional inverse design techniques, such as the Feedforward Inverse Model (FIM), offering improved generalization and prediction accuracy. The approach begins with a FIM to generate initial design estimates, followed by an iterative correction strategy inspired by the Hybrid Inverse-Forward Residual Refinement Network (HiFR2-Net), which we call IRC-Net. Experiments demonstrate that the IRC-Net achieves substantial improvements in prediction accuracy compared to traditional single-stage networks, validated through statistical metrics, full-wave electromagnetic simulations, and measurements. To validate the proposed framework, we first design and fabricate a three-resonance SIW structure. Next, we apply the trained IRC-Net model to predict the geometry of a four-resonance structure based on its desired frequency response. Both designs are fabricated and tested, showing strong agreement between the simulated, predicted, and measured results, confirming the effectiveness and practicality of the proposed method.
Keywords Inverse Design  $\cdot$ Iterative Residual Correction  $\cdot$ Deep Learning  $\cdot$ Substrate Integrated Waveguide (SIW)  $\cdot$ Multimode Resonators  $\cdot$ Residual Neural Networks  $\cdot$ Neural Networks  $\cdot$ Inverse Electromagnetic Modeling  $\cdot$ Smart Filter Design
@@ 1 Introduction
The Ku-band plays a pivotal role in modern communication systems, supporting critical applications such as satellite links, radar systems, and point-to-point wireless communications. Offering larger bandwidth allocations compared to lower-frequency bands, the Ku-band has become essential for contemporary telecommunication infrastructures. The rising demand for high-capacity, data-intensive applications, including HD video streaming, real-time data transmission, and broadband satellite Internet, has further intensified the need for optimized Ku-band systems, especially with the proliferation of low Earth orbit (LEO) and geostationary satellite networks [1, 2, 3].
Despite its advantages, Ku-band communication faces several challenges. Signals at higher frequencies are more prone to attenuation, interference, and environmental factors, such as rain fading, which limit the effective transmission range [4, 5]. As a result, there is a critical need for highly efficient microwave components that meet stringent performance requirements. Among these components, microwave filters play a crucial role in suppressing unwanted frequencies while preserving the integrity of the signal. The precise design of these filters is essential to ensure effective channelization and maintain communication quality[6, 7, 8].
SIW technology has emerged as a revolutionary solution for high frequency applications, particularly in the Ku-band. By embedding rows of conductive vias into a dielectric substrate, SIW combines the advantages of conventional waveguides with the ease and cost-efficiency of planar circuit fabrication. This approach enables compact, low-loss and mass producible components, suitable for integration with antennas, amplifiers, and other microwave circuits [9, 10].
In general, designing various electromagnetic structures can sometimes be challenging. However, if one can intelligently identify the resonances of a structure and effectively control them, it becomes possible to design different functional devices, such as filters, by appropriately shaping the structures [11, 12, 13].
@@
Traditional design approaches for electromagnetic structures, such as antenna arrays and microwave components, often rely on iterative full-wave simulations, heuristic optimization, or expert-driven trial-and-error processes. These methods, while accurate, are computationally intensive and time-consuming, particularly for high-dimensional and nonlinear design spaces. In recent years, data-driven approaches—particularly deep learning (DL)—have emerged as powerful alternatives. The DL has shown promise in both forward modeling and inverse design. Numerous studies have demonstrated the effectiveness of DL-based inverse design in a wide range of electromagnetic applications, including phase retrieval in phase antenna arrays [14], prediction of pulse widths and delays in time-modulated antenna arrays to for synthesizing of multi-beams [15], automatic metasurface synthesis over broad frequency bands [16, 17], reconfigurable terahertz antenna design [18], and electromagnetic inverse scattering [19, 20].
Building on recent advancements, researchers have increasingly applied deep learning techniques to both the forward modeling and inverse design of high-frequency microwave components. In [21], a single-stage neural network utilizing smooth ReLU and sigmoid activations is proposed for inverse modeling and parameter extraction of microwave filters. However, the model exhibits signs of overfitting, is trained on only 41 frequency samples for S-parameters, and lacks experimental validation, limiting its robustness and practical applicability.
Most research focus exclusively on either the forward or inverse problem, missing the opportunity to enforce physics consistency via a bi-directional training loop. In [22], an R-DBN-based inverse model is proposed for the first time for modeling of a sixth-order and a fourth-order multicoupled cavity filter. The learning time is relatively long, no physical implementation is reported, and only 30 S-parameter samples are used, which raises concerns about generalizability. Likewise, [23], an attention-based deep neural network combining convolutional neural network, joint channel–spatial attention, and residual connections has been proposed for forward modeling of spatio-temporal modulated (STM) non-reciprocal microwave circuits, aiming to improve prediction accuracy and generalization across various circuit configurations.
@@
Among the limited work focused on SIW filters, [24] presents a method combining an inverse neural network with a K-impedance inverter model to synthesize a fifth-order H-plane SIW iris filter. This approach utilizes an equivalent de-embedding technique to directly extract the generalized scattering matrix, thereby avoiding conversion errors associated with equivalent rectangular waveguide models. Although results show good agreement with full-wave simulations, the method remains constrained to a fixed-order filter, is trained on a relatively small dataset, and does not accommodate a range of SIW topologies or structural variations. Furthermore, it employs conventional neural network architectures, lacking recent advances in deep learning design and training methodologies.
To address these limitations, this paper proposes a novel three-stage deep learning framework for the inverse modeling of SIW multimode resonators. The framework begins with a Feedforward Inverse Model (FIM) that generates an initial estimate of the resonator design parameters. This estimate is then refined by a Hybrid Inverse-Forward Residual Refinement Network (HiFR2-Net), which combines forward modeling with residual correction to enhance prediction accuracy. Finally, an Iterative Residual Correction Network (IRC-Net) further reduces residual errors through multiple refinement stages.
By integrating coarse-to-fine prediction, forward-consistency enforcement, and iterative residual learning, the proposed framework enables high-precision inverse design without relying on extensive electromagnetic simulations. Experimental validations confirm the method’s effectiveness, demonstrating low prediction error and strong generalization performance across diverse design scenarios.
The remainder of this paper is organized as follows: Section 2 presents the design principles of substrate integrated waveguide (SIW) multimode resonator structures. A proposed configuration supporting multiple resonant modes within the Ku-band is introduced, and a comprehensive parametric study is conducted to evaluate the influence of key structural parameters on the resonance characteristics—crucial for compact and tunable filter designs. Section 3 describes the proposed deep learning framework for inverse modeling, including dataset preparation, model architecture, and training strategy. Section 4 provides a detailed analysis of the model’s learning behavior, generalization performance, and simulation-based validation. Finally, Section 5 concludes the paper with a summary of findings and suggestions for future research directions.
@@ 2.1 SIW Structure Overview
SIW typically consists of a dielectric substrate sandwiched between two metal planes, with rows of metallized via holes forming the sidewalls of the waveguide. These vias confine the electromagnetic fields within the substrate, supporting the propagation of transverse electric (TE) modes.
The effective width ($W_{eff}$), which determines its cutoff frequency, can be approximated by:
@@
$p<\frac{\lambda_{g}}{5},d<2p$ (2)
Here, $\lambda_{g}$ is the guided wavelength within the SIW. Adhering to these constraints helps prevent leakage between the vias and maintains the integrity of the waveguide mode.
Multi-resonant SIW structures are realized by placing metallic posts within the SIW waveguide. These posts exhibit inductive behavior, and their proper placement enables the creation of multiple resonant frequencies. By carefully selecting the number, size, and positions of the posts, the resonances can be brought closer together, resulting in a bandpass filtering response [25, 26]. The primary degrees of freedom in the design process are the spacing between the posts, the radius of each post, and their specific locations within the structure.
Our main objective is to demonstrate the occurrence of multiple resonances and analyze how these resonances vary with different structural parameters. Therefore, we choose design parameters that yield multiple resonances at distinct frequencies. Since the posts introduce inductive loading, they are strategically placed at locations with maximum current density. Given that the field distribution in SIW structures is analogous to that of rectangular waveguides, the high-current regions can be identified with relative ease, allowing for an effective and informed post placement strategy.
In the following, we proceed to design an SIW structure incorporating the aforementioned characteristics. The design aims to achieve a multi-resonant response by exploiting the inductive nature of the posts and carefully tuning their geometrical and positional parameters. Through full-wave simulations, we analyze how variations in these parameters influence the resonant frequencies and demonstrate the potential of such structures for compact, planar filter implementations.
@@ 2.2 Physical Layout, Parameter Definition and Result of Simulation
As mentioned earlier, the posts act as inductive elements and should therefore be positioned at locations of maximum current within the SIW structure. It is also well established that increasing the post diameter results in higher inductance. In conventional bandpass filter design, these inductive elements are typically arranged in ascending order—i.e., from smaller to larger diameters—to achieve the desired frequency response.
However, in this study, the objective is to design a multi-mode resonator structure with widely spaced resonant frequencies, allowing us to examine how geometric variations influence the frequency shift of each individual resonance. To achieve this, the posts are deliberately arranged in descending order of diameter, from larger to smaller, to increase the separation between resonant modes.
@@
The design aims to achieve three distinct resonant frequencies within the operating band. It is also possible to obtain additional resonances within the band by increasing the number of resonators and the overall length of the structure. Additionally, multi-section structures can be employed to introduce varying impedance structures, offering greater degrees of freedom for tuning the positions of the resonant frequencies, either bringing them closer together or spacing them further apart. However, the primary objective here is to demonstrate that the proposed structure is capable of generating designer-specified resonances. The geometrical parameters for the three in-band resonances are summarized in Table 1. The values were initially selected on the basis of established design guidelines.
To evaluate the electromagnetic performance of the proposed filter, full-wave simulations were performed using CST Microwave Studio. Figure 3 presents the simulated S parameters ($S_{11}$ and $S_{21}$) for the three in band resonances. The reflection coefficient ($S_{11}$) exhibits sharp notches at the desired center frequencies, indicating strong impedance matching. Meanwhile, the transmission coefficient ($S_{21}$) shows a low insertion loss throughout the passband, confirming efficient signal propagation and validating the effectiveness of the designed SIW multimode resonator.
Table 1: Geometrical parameters of the multi-mode SIW structure for three in-band resonances.
@@
$R_{1}$ Radius of the first coupling via 0.25
$R_{2}$ Radius of the second coupling via 0.8
$R_{3}$ Radius of the third coupling via 0.9
$D_{1}$ Distance between the first and second posts 5.8
@@
$d$ Diameter of the sidewall vias forming the SIW boundary 0.8
$p$ Center-to-center pitch between adjacent sidewall vias 1.3
@@
Figure 3: Simulated S parameters ($S_{11}$ and $S_{21}$) of the proposed multi-mode SIW structure with 3 in-band resonances.

### v2

@@ Abstract
Designing high-performance substrate-integrated waveguide (SIW) filters with both closely spaced and widely separated resonances is challenging. Consequently, there is a growing need for robust methods that reduce reliance on time-consuming electromagnetic (EM) simulations.
In this study, a deep learning-based framework was developed and validated for the inverse design of multi-mode SIW filters with both closely spaced and widely separated resonances. A series of SIW filters were designed, fabricated, and experimentally evaluated. A three-stage deep learning framework was implemented, consisting of a Feedforward Inverse Model (FIM), a Hybrid Inverse-Forward Residual Refinement Network (HiFR2-Net), and an Iterative Residual Correction Network (IRC-Net).
The design methodology and performance of each model were systematically analyzed. Notably, IRC-Net outperformed both FIM and HiFR2-Net, achieving systematic error reduction over five correction iterations. Experimental results showed a reduction in mean squared error (MSE) from 0.00191 to 0.00146 and mean absolute error (MAE) from 0.0262 to 0.0209, indicating improved accuracy and convergence.
The proposed framework demonstrates the capability to enable robust, accurate, and generalizable inverse design of complex microwave filters with minimal simulation cost. This approach is expected to facilitate rapid prototyping of advanced filter designs and could extend to other high-frequency components in microwave and millimeter-wave technologies.
keywords:
Deep Learning , Inverse Design , Iterative Residual Correction , Substrate Integrated Waveguide (SIW) Filter , Multimode Resonators , Ku-band
journal: International Journal of Electronics and Communications
Accepted Manuscript Notice
This is the accepted version of the manuscript published in:
AEU - International Journal of Electronics and Communications
DOI: https://doi.org/10.1016/j.aeue.2025.156003
© 2025 Elsevier Ltd. All rights reserved.
[…]
[inst1] organization=Department of Electrical Engineering, Iran University of Science and Technology (IUST), city=Tehran, country=Iran
[…]
[inst2] organization=Department of Electrical Engineering, Amirkabir University of Technology (Tehran Polytechnic), city=Tehran, country=Iran
[…]
[inst3] organization=Department of Electrical and Computer Engineering, Isfahan University of Technology (IUT), city=Isfahan, country=Iran
[…]
[inst4] organization=Department of Electrical and Computer Engineering, Concordia University, city=Montreal, country=Canada
{graphicalabstract}[Uncaptioned image]
{highlights}
Novel SIW filters with both close and wideband resonances are realized
Deep learning trained on large SIW dataset enables accurate inverse design
HiFR2-Net unifies inverse, forward, and residual learning for refinement
IRC-Net applies iterative correction to reduce design estimation error
Experimental results confirm high precision and strong real-world performance
@@ 1 Introduction
Resonant microwave structures underpin many modern filter designs, offering high quality-factor responses through carefully engineered cavities and resonators. Microwave filters play a crucial role in suppressing unwanted frequencies while preserving the integrity of the signal. The precise design of these filters is essential to ensure effective channelization and maintain communication quality [1, 2, 3]. Among the various classes of filters, multimode filters are particularly significant due to their compact size and ability to support multiple frequency bands, meeting the requirements and demands of modern communication systems, especially multi-standard wireless communication[4, 5].
Multimode microwave filters have been designed and implemented across a range of frequency bands [6]. However, the Ku-band stands out due to its growing importance in supporting high-capacity, data-intensive applications such as HD video streaming, real-time data transmission, and broadband satellite Internet. The growing use of LEO and geostationary satellites increases the need for high-performance components in this frequency range. To address demands like compact size, integration, and high selectivity, various designs have been proposed. For example, [7] presented a planar bandpass filter using stepped-impedance resonators and multiple coupling paths. It offers a good balance between size, performance, and cost for PCB-based systems. In another study, [8] analyzed the Starlink Ku-band downlink, highlighting the importance of filters in ensuring stable broadband in dense satellite networks. Consequently, the development of compact and efficient multimode filters operating in the Ku-band has become a critical objective in the design of next-generation communication systems.
While multimode resonant structures have been implemented using technologies such as microstrip [9] and conventional waveguides [10], SIW technology offers a compelling alternative for high-frequency applications, particularly in the Ku-band. By embedding rows of conductive vias into a dielectric substrate, SIW preserves the low loss and high quality factor of traditional waveguides while enabling planar fabrication, compact size, and cost-effective integration [11]. These features make SIW well-suited for modern microwave systems, allowing seamless integration with antennas, amplifiers, and other RF components. In this context, [12] investigated interconnect solutions for SIW structures at millimeter-wave frequencies, addressing key challenges in compact integration. Additionally, [13] reviewed recent progress in SIW-based antenna technologies, highlighting the growing role of reconfigurable designs and AI-assisted methods in advancing 5G and MIMO systems. Together, these efforts demonstrate SIW’s potential as an effective platform for compact, efficient, and high-performance multimode resonators used in Ku-band single- and multi-band microwave filters.
@@
Traditional design methods for electromagnetic structures—including antenna arrays and microwave components—typically depend on iterative full-wave simulations, heuristic optimizations, and labor-intensive, expert-driven trial-and-error approaches. While accurate, these methods are computationally intensive and time-consuming, particularly for high-dimensional and nonlinear design spaces. In recent years, data-driven approaches, particularly deep learning (DL), have emerged as powerful alternatives. The DL has shown promise in both forward modeling and inverse design. Numerous studies have demonstrated the effectiveness of DL-based methods in a wide range of electromagnetic applications, including phase retrieval in phased antenna arrays [14], pulse width and delay prediction in time-modulated antenna arrays for multibeam synthesis [15], automatic metasurface synthesis over broad frequency bands [16, 17], inverse design of reconfigurable and multiband antennas using deep learning frameworks [18, 19], electromagnetic inverse scattering [20], microwave biomedical diagnostics using S-parameter profiles instead of imaging data [21], and fast and accurate surrogate modeling of antennas using convolutional neural networks to predict near- and far-field parameters with significant speedup over full-wave simulations [22].
Deep learning approaches to microwave component design generally fall into two categories: Forward modeling and inverse design. Forward modeling networks take device geometry as input and predict its scattering response (e.g., S-parameters). In [23], a deep neural network (DNN) is employed to predict the resonant frequency of E-shaped patch antennas (ESPAs) with high accuracy. The model is trained on a dataset of 144 simulated ESPAs and validated experimentally, demonstrating the potential of DNNs for rapid antenna analysis. Similarly, [24] introduces an attention-based deep neural network combining convolutional neural network, joint channel–spatial attention, and residual connections that have been proposed for forward modeling of spatio-temporal modulated (STM) non-reciprocal microwave circuits, aiming to improve prediction accuracy and generalization across various circuit configurations.
Inverse design methods, in contrast, aim to infer a geometry for a given desired response. Early examples include Kabir et al., who introduced neural-network inverse modeling for microwave filters, formulating the problem where the inputs are desired electrical specifications and the outputs are filter dimensions [25]. Since then, researchers have tried to overcome implementation challenges and reduce errors in model parameter extraction. In [26], a single-stage neural network utilizing smooth ReLU and sigmoid activations is proposed for inverse modeling and parameter extraction of microwave filters. However, the model exhibits signs of overfitting, is trained on only 41 frequency samples for S-parameters, and lacks experimental validation, limiting its robustness and practical applicability. In [27], an R-DBN-based inverse model is proposed for the first time for modeling a sixth-order and a fourth-order multicoupled cavity filter. The learning time is relatively long, no physical implementation is reported, and only 30 S-parameter samples are used, which raises concerns about generalizability. Recent advances, such as [28]’s generalized deep-learning inverse design for complex multiport RF components, enable rapid synthesis of arbitrary-shaped electromagnetic structures with tailored radiative/scattering properties. As demonstrated in [29], a DL-assisted optimization method is introduced to facilitate filter design; however, it still relies on PSO optimization and full-wave EM simulations, and thus is not a purely data-driven approach.
@@
Among the limited work focused on SIW filters, [30] presents a method combining an inverse neural network with a K-impedance inverter model to synthesize a fifth-order H-plane SIW iris filter. While this method avoids conversion errors via de-embedding, thereby avoiding conversion errors associated with equivalent rectangular waveguide models, it is limited to a fixed-order filter, is trained on a relatively small dataset, and does not accommodate a range of SIW topologies or structural variations. Furthermore, it employs conventional neural network architectures, lacking recent advances in deep learning design and training methodologies. Likewise, [31] presents an invertible neural networks (INNs) for designing D-band SIW filters. Although this method addresses non-unique solutions, it covers only two design parameters and is trained on a small dataset. Similarly, [32] achieves a 24% fractional bandwidth in Ku-band SIW filters with the help of optimization using a feed-forward backpropagation neural network. However, their approach relies on a simple ANN trained on a limited dataset, which constrains modeling accuracy and generalizability. In addition to ANN-based methods, [33] presents a binary optimization approach with notable speedup, yet the limitations in adaptability and generalization remain unaddressed. To better contextualize the novelty of this work, a comparative table summarizing recent SIW inverse design techniques, including their datasets, methods, and key results, is provided in Table 1.
Table 1
Comparative overview of inverse design methods
Reference Method Dataset Size Complexity Generalizability Runtime Cost
[30] ANN 100 Low Moderate Fast (few ms)
[31] ANN 150 Low Moderate Fast (few ms)
[32] ANN 386 Low Moderate Fast (few ms)
[33] Genetic Algorithm - Medium Low Slow (5–10 s)
This paper presents the design, fabrication, and experimental validation of multimode SIW filters capable of generating multiple resonances, including both closely spaced and widely separated frequency responses. This outcome is enabled by a novel deep learning framework structured into three progressive stages. This achievement is driven by a novel deep learning framework structured into three progressive stages. First, a conventional Feedforward Inverse Model (FIM) is employed to generate an initial estimate of the design parameters. This model establishes a baseline prediction by mapping S-parameters to the corresponding design parameters. Next, this initial estimate is enhanced using a Hybrid Inverse-Forward Residual Refinement Network (HiFR2-Net), which combines inverse prediction with forward validation and residual correction. By incorporating the forward response within the learning process, this hybrid approach refines the design parameter estimation through iterative residual feedback, improving accuracy and robustness. To further optimize the prediction process, an Iterative Residual Correction Network (IRC-Net) is introduced. This architecture integrates principles from classical iterative refinement algorithms [34, 35] into the deep learning framework. Unlike the previous models, the IRC-Net directly refines the initial design estimate through multiple residual correction cycles, without reprocessing the scattering parameters. This mathematical integration, inspired by IRC methods originally developed for solving Volterra integral equations and linear systems, allows for systematic error reduction and improved convergence. Compared to the baseline FIM and the hybrid HiFR2-Net, our proposed IRC-Net demonstrates superior performance in inverse design tasks, enabling high-precision parameter estimation with minimal reliance on extensive electromagnetic simulations. Experimental validation confirms the effectiveness of this approach, showing low prediction error and strong generalization capabilities across diverse design scenarios.
To provide a comprehensive overview of our methodology and results, the remainder of this paper is organized as follows. Section II presents the design of the multimode resonant SIW structure, detailing its physical layout, parameter definitions, and the results of simulation-based investigations. Section III introduces our deep learning methodology, comprising the preparation of a simulation-driven dataset, development of a feedforward inverse model, and refinement through hybrid inverse-forward and iterative residual correction networks. Section IV discusses the learning behavior and performance evaluation of the proposed models, including statistical analysis and practical validation through simulation and fabrication. Finally, Section V concludes the paper, summarizing key findings and outlining directions for future research.
@@ 2.1 SIW Structure Overview
SIW typically consists of a dielectric substrate sandwiched between two metal planes, with rows of metalized via holes forming the sidewalls of the waveguide. These vias confine the electromagnetic fields within the substrate, supporting the propagation of transverse electric (TE) modes.
The effective width ($W_{eff}$), which governs the cutoff frequency of the structure, can be approximated as follows [36]:
@@
$p<\frac{\lambda_{g}}{4},d<p$ (2)
here, $\lambda_{g}$ is the guided wavelength within the SIW. Adherence to these constraints helps prevent leakage between the vias and maintains the integrity of the waveguide mode. See [37] for more details.
Multi-resonant SIW structures are formed by embedding metallic posts within the SIW waveguide. These posts exhibit inductive behavior, and their proper placement enables the creation of multiple resonant frequencies. By carefully selecting the number, size, and positions of the posts, the resonances can be brought closer together, resulting in a bandpass filtering response [38, 39]. Furthermore, by increasing the frequency spacing between the resonances, it is also possible to design dual-band or multi-band filters, providing more flexibility for various communication and sensing applications [40, 41, 42]. The primary degrees of freedom in the design process are the spacing between the posts, the radius of each post, and their specific locations within the structure.
Our main objective in this section is to demonstrate the occurrence of multiple resonances and to analyze how these resonances vary with different structural parameters. Accordingly, design parameters are selected to produce multiple resonances at distinct frequencies. In the case of filter design, where closely spaced resonances are desired, the inductive posts are placed at locations with maximum current density, since they exhibit inductive behavior. Given that the field distribution in SIW structures is analogous to that of rectangular waveguides, these high-current regions can be easily identified, enabling an effective and well-informed post placement strategy.
The following section presents the design of an SIW structure incorporating the aforementioned characteristics. The design aims to achieve a multi-resonant response by exploiting the inductive nature of the posts and carefully tuning their geometrical and positional parameters. Full-wave simulations are employed to analyze how variations in these parameters affect the resonant frequencies, thereby demonstrating the potential of such structures for compact, planar filter implementations.
@@ 2.2 Physical Layout, Parameter Definition and Result of Simulation
As mentioned earlier, the posts act as inductive elements and should therefore be positioned at locations of maximum current within the SIW structure. It is also well established that increasing the post diameter results in higher inductance. In conventional bandpass filter design, these inductive elements are typically arranged in ascending order—i.e., from smaller to larger diameters to achieve the desired frequency response.
In this section, the focus is on designing a multi-mode resonator structure that enables the investigation of how geometric variations influence the frequency shift of each individual resonance. To achieve this, the posts are arranged in descending order of diameter, from larger to smaller, which facilitates differentiation among the resonant modes and allows for observation of their spacing, whether close or wide apart.
@@
This structure is capable of generating four distinct resonant frequencies within the operating band. Moreover, by bringing two of the resonances closer together, it is possible to realize structures with three or even two resonant frequencies, offering more flexibility for various filter design requirements. It is also possible to obtain additional resonances within the band by increasing the number of resonators and the overall length of the structure. Additionally, multi-section structures can be used to create different impedance profiles, providing more flexibility in tuning the resonant frequencies. This allows the designer to either bring the resonances closer together or space them further apart, depending on the desired filter response. However, the main goal here is to demonstrate that the proposed structure can generate resonances at frequencies specified by the designer. The geometric parameters used to achieve the four in-band resonances are listed in Table 2. These values were initially chosen based on well-known design guidelines.
To evaluate the electromagnetic performance of the proposed filter, full-wave simulations were performed using CST Microwave Studio. Figure 3 presents the simulated S parameters ($S_{11}$ and $S_{21}$) for the four in-band resonances. The reflection coefficient ($S_{11}$) exhibits sharp notches at the desired center frequencies, indicating strong impedance matching. Meanwhile, the transmission coefficient ($S_{21}$) shows a low insertion loss throughout the passband, confirming efficient signal propagation and validating the effectiveness of the designed SIW multimode resonator.
Table 2
Geometrical parameters of the multi-mode SIW structure for three in-band resonances.
@@
$R_{1}$ Radius of the first coupling via 0.2
$R_{2}$ Radius of the second coupling via 0.4
$R_{3}$ Radius of the third coupling via 0.8
$D_{1}$ Distance between the first and second posts 5.5
@@
$d$ Diameter of the sidewall vias 0.8
$p$ Spacing between sidewall vias 1.3
@@
Figure 3: Simulated S parameters ($S_{11}$ and $S_{21}$) of the proposed multi-mode SIW structure with 4 in-band resonances.